OpenStudy (anonymous):

(x + 1)(x - 1) factoring trinomial

OpenStudy (anonymous):

foil method!

OpenStudy (anonymous):

Use the Foil method F = Multiply First set of numbers O = Multiply Outer set of numbers I = Multiply Inner set of numbers L = Multiply Last set of numbers F = x * x =x^2 O = x * -1 = -1x I = x *1 = 1x L = 1 * -1 = -1 Put it together